# REVS: Unlearning Sensitive Information in Language Models via Rank Editing in the Vocabulary Space

### v2

@@ Appendix F Dataset Curation
While datasets such as Who’s Harry Potter (Eldan and Russinovich, 2023), WMDP (Li et al., 2024), and TOFU (Maini et al., 2024) are widely used, they focus on concept unlearning. Concept unlearning differs fundamentally from the challenge of unlearning text-based sensitive information, where the goal is to obfuscate a single piece of sensitive information rather than broad knowledge of a concept, rendering these datasets unsuitable for our specific goals.
To the best of our knowledge, no publicly available dataset contains sensitive information that models naturally memorize. To address this gap, we curated a novel English language dataset specifically designed to evaluate unlearning methods on real-world, organically memorized email addresses for each model. To provide a more comprehensive and diverse evaluation, we also curated a synthetic English language dataset in which the sensitive token sequences are numbers (as opposed to the character-based sequences in the email dataset).
F.1 Email Dataset
To create the Email dataset, we identified sentences containing email addresses using Presidio (Microsoft, 2024). We then prompted the model with the prefix text preceding the email address and verified that it generated the exact email within the next 20 tokens using greedy decoding. The target token sequence corresponds to the string before @email.com. This dataset enables evaluation of unlearning methods on data that the model has naturally memorized from its training corpus. We show sample sentences from the email dataset with the actual emails obfuscated in Table 11.

### v3

@@ Appendix F Dataset Curation
While datasets such as Who’s Harry Potter (Eldan and Russinovich, 2023), WMDP (Li et al., 2024), and TOFU (Maini et al., 2024) are widely used, they focus on concept unlearning. Concept unlearning differs fundamentally from the challenge of unlearning sensitive information, where the goal is to obfuscate a single piece of sensitive information rather than broad knowledge of a concept, rendering these datasets unsuitable for our specific goals.
To the best of our knowledge, no publicly available dataset contains sensitive information that models naturally memorize. To address this gap, we curated a novel English language dataset specifically designed to evaluate unlearning methods on real-world, naturally memorized for each model. To provide a more comprehensive and diverse evaluation, we also curated a synthetic English language dataset in which the sensitive token sequences are numbers (as opposed to the character-based sequences in the email and URL datasets).
F.1 Naturally Memorized Dataset
Curating a dataset of naturally memorized sensitive information involves identifying sensitive instances in the model’s training data and verifying their memorization. However, determining both the sensitivity of an instance and its memorization is computationally expensive and requires specialized tools like Presidio (Microsoft, 2024).
Existing work, such as Elazar et al. , use regex-based filtering on datasets like The Pile (Gao et al., 2020) to identify sensitive information. However, these methods produce many false positives—instances flagged as sensitive but that are not genuinely so.
To address this, we leveraged the “Training Data Extraction Challenge” [^2] https://github.com/google-research/lm-extraction-benchmark/blob/master/detailed_description.pdf, a subset of The Pile known to contain sensitive information. This subset comprises 15,000 sentences in which GPT Neo exhibited verbatim memorization of sentence suffixes given their prefixes. We used this subset as a strong candidate for naturally memorized sensitive data.
We analyzed the dataset using Presidio (Microsoft, 2024) to identify various types of personally identifiable information (PII). For each identified PII instance, we tested whether the model could reproduce it exactly when given the original context prefix. The distribution of identified and memorized PII types by Llama 3 8B is summarized in Table 10.
PII Type Identified Memorized
URLs 17,233 203
Email addresses 1,631 205
Phone numbers 357 0
IP addresses 55 0
Table 10: Distribution of identified and memorized PII types by Llama 3 8B in the dataset
F.1.1 Emails Dataset
To curate the email dataset, we used Presidio to identify sentences containing email addresses. We then test memorization by providing the model with the text preceding each email address and checking whether it could generate the exact email within the next 50 tokens using greedy decoding. The target token sequence for unlearning is defined as the substring preceding “@email.com.”
Our results show that Llama 3 8B memorized 205 email addresses from the dataset, while GPT J 6B memorized 288. Sample sentences from this dataset, with emails obfuscated, are presented in Table 11.
F.1.2 URLs Dataset
For the URL dataset, we identified sentences containing URLs using Presidio. We filtered out URLs shorter than 40 characters to exclude generic addresses (e.g., “https://www.gnu.org”) and URLs containing the word “License.” Our analysis found that Llama 3 8B memorized 203 URLs meeting these criteria. We did not evaluate URL memorization for GPT J 6B.